# Generalization under Byzantine & Poisoning Attacks: Tight Stability Bounds in Robust Distributed Learning

### v1

@@ Appendix F Refined bounds under bounded heterogeneity and bounded variance assumptions
Consider the setting described in Section 2, under either Byzantine failure or poisoning attacks. Let $\mathcal{A}$ an $\varepsilon$-uniformly stable. Assume $\forall z\in\mathcal{Z},\ell(\cdot,z)\leq\ell_{\infty}$, then we have that for any $10<\delta<$ 1, with probability at least $1-\delta$:
$|R_{\mathcal{H}}(\mathcal{A}(\mathcal{S}))-\widehat{R}_{\mathcal{H}}(\mathcal{A}(\mathcal{S}))|\in\mathcal{O}\Bigg{(}\min\bigg{\{}\varepsilon\ln(|\mathcal{H}|m)\ln(\frac{1}{\delta})+\frac{\ell_{\infty}}{\sqrt{|\mathcal{H}|m}}\sqrt{\ln(\frac{1}{\delta})},\\ \left(\sqrt{\varepsilon\ell_{\infty}}+\frac{\ell_{\infty}}{\sqrt{|\mathcal{H}|m}}\right)\sqrt{\ln(\frac{1}{\delta})}\bigg{\}}\Bigg{)}$
Poisoning attacks.
By adapting the optimization bound technique from [15, Theorem 4] to $\mathrm{GD}$ algorithm with $\mathrm{SMEA}$ under poisoning attacks, we show that for a constant number of steps—specifically, when the step count satisfies $T\geq\ln(4\frac{f+1/m}{n-f}+\kappa)/\ln(1-\frac{\mu}{L})\in\mathcal{O}(1)$—the expected excess risk is of the order of:
@@
Additionally, we provide a high-probability bound on the excess risk using Lemma G.1. Specifically, for any $10<\delta<$ 1, with probability at least $1-\delta$ the excess risk lies within:
@@
In the absence of a minimax statistical lower bound for the heterogeneous case—and acknowledging that the comparison is imperfect due to differing assumptions—it remains informative to compare our resulting bound with those in [47, 48], as well as with Proposition 1 in [4].
Byzantine failure attacks.
@@
Under Byzantine failure attacks, the derivation proceeds similarly and yields an analogous bound, with a dependence on $\sqrt{\kappa}$ replacing the $\frac{f}{n-f}$ term in the uniform stability bound.

### v2

@@ Appendix G HIGH-PROBABILITY EXCESS RISK BOUNDS FOR ROBUST DISTRIBUTED LEARNING
Consider the setting described in Section 2, under either Byzantine failures or data poisoning. Let $\mathcal{A}$ an $\varepsilon$-uniformly stable. Assume $\forall z\in\mathcal{Z},\ell(\cdot,z)\leq\ell_{\infty}$, then we have that for any $0<\delta<1$, with probability at least $1-\delta$:
$|R_{\mathcal{H}}(\mathcal{A}(\mathcal{S}))-\widehat{R}_{\mathcal{H}}(\mathcal{A}(\mathcal{S}))|\in\mathcal{O}\Bigg(\min\bigg\{\varepsilon\ln(|\mathcal{H}|m)\ln(\frac{1}{\delta})+\frac{\ell_{\infty}}{\sqrt{|\mathcal{H}|m}}\sqrt{\ln(\frac{1}{\delta})},\left(\sqrt{\varepsilon\ell_{\infty}}+\frac{\ell_{\infty}}{\sqrt{|\mathcal{H}|m}}\right)\sqrt{\ln(\frac{1}{\delta})}\bigg\}\Bigg)$
Data poisoning.
By adapting the optimization bound technique from Farhadkhani et al. (2024b, Theorem 4) to $\mathrm{GD}$ algorithm with $\mathrm{SMEA}$ under data poisoning, we show that for a constant number of steps—specifically, when the step count satisfies $T\geq|\ln(4\frac{f+1/m}{n-f}+\kappa)/\ln(1-\frac{\mu}{L})|\in\mathcal{O}(1)$—the expected excess risk is of the order of:
@@
Additionally, we provide a high-probability bound on the excess risk using Lemma G.1. Specifically, for any $0<\delta<1$, with probability at least $1-\delta$ the excess risk lies within:
@@
In the absence of a minimax statistical lower bound for the heterogeneous case—and acknowledging that the comparison is imperfect due to differing assumptions—it remains informative to compare our resulting bound with those from Yin et al. (2018); Zhu et al. (2023), as well as Allouah et al. (2023a, Proposition 1).
Byzantine failures.
@@
Under Byzantine failures, the derivation proceeds similarly and yields an analogous bound, with a dependence on $\sqrt{\kappa}$ replacing the $\frac{f}{n-f}$ term in the uniform stability bound.
Appendix H ADDITIONAL DISCUSSIONS
Why the difference in threat models (Byzantine failures versus data poisoning) affects generalization but not optimization?
The key intuition behind why the difference in threat models affects generalization but not optimization is as follows. Optimization error analysis focuses on how accurately the true descent direction is estimated at each step, whereas generalization error analysis examines the algorithm’s sensitivity to individual training samples, independent of how informative the estimated direction is for optimization. In the worst case, both types of corruption similarly impair the estimation of the optimization direction. However, stability analysis benefits from the additional regularity present in data poisoning, where corrupted gradients still correspond to gradients of a smooth loss function (i.e., they remain Lipschitz)—unlike Byzantine failures, which allow for arbitrary, potentially non-smooth corruptions.
Specifically, optimization error analysis relies on the smoothness of the empirical loss over honest workers (e.g., Theorem 1 in Allouah et al., 2023a) to control its decrease along the optimization trajectory, typically via terms like $\widehat{R}_{\mathcal{H}}(\theta_{t+1})-\widehat{R}_{\mathcal{H}}(\theta_{t})$ or $\|\nabla\widehat{R}_{\mathcal{H}}(\theta_{t})\|^{2}_{2}$. This smoothness-based control applies equally to both threat models, as the Lipschitzness of corrupted gradients is not invoked, and yields optimal bounds. In contrast, stability analysis aims to bound the difference $\ell(\theta_{t},z)-\ell(\theta^{\prime}_{t},z)$ between the losses along two optimization trajectories $\theta_{t}$ and $\theta^{\prime}_{t}$ produced by neighboring datasets. Here, smoothness can be exploited under data poisoning—where corrupted updates remain gradients of a smooth loss function (cf. proof of Theorem 3.3)—but not under Byzantine failures, which require an intermediate comparison step (cf. proof sketch in Section C.2).
Furthermore, prior work by Allouah et al. (2023a) have shown that for the optimization error, $(f,\kappa)$-robustness with $\kappa\in\mathcal{O}(f/n)$ suffices to achieve a lower bound obtained with poisoned data. A matching upper-bound is obtained for both Byzantine failures and data poisoning threat models (Allouah et al., 2023a, Theorem 1 & Proposition 1), meaning that we cannot hope to improve it, even using the Lipschitzness of corrupted gradients under data poisoning.
On the relevance of worst-case bounds to practical learning scenarios.
Although variations in attack strength and implementation may contribute to the observed gap in generalization performance between the two threat models, our results indicate that this gap cannot be explained solely by the fact that data poisoning attacks used in practice are often empirically weak (suboptimal within their threat model), whereas Byzantine attacks used in practice tend to be empirically strong (near-optimal). Rather, the gap also reflects an inherent difference in harm capacity (Section 4).
We recall that a gap in test accuracy has been consistently reported in prior work on realistic datasets, further motivating our investigation. For example, (Allouah et al., 2023a, Table 2) compares Byzantine attacks such as $\mathrm{ALIE}$, $\mathrm{FOE}$, and $\mathrm{SF}$ with data poisoning attacks such as $\mathrm{LF}$ and $\mathrm{Mimic}$, showing a clear disparity in their impact.
Does our conclusions would fundamentally change for other popular first-order optimization methods (e.g. momentum, variance reduction, multiple local steps)?
We believe our proof techniques are broadly applicable to robust first-order optimization, so the fundamental results should remain consistent. Quantitatively characterizing this phenomenon for specific algorithms is an interesting direction for future work. More concretely, similar recursions can be derived for stochastic heavy-ball momentum, using the proof techniques of Section C.2 (under Byzantine failures), and the refined approach in the proof of Theorem 3.3 (under data poisoning). These derivations reveal a similar gap as in robust distributed $\mathrm{GD}$ and $\mathrm{SGD}$: a $\sqrt{\kappa}$ dependency appears under Byzantine failures, while the tighter bound $\frac{f}{n-f}$ arises under data poisoning. However, fully analyzing these more complex algorithms introduces additional technical challenges. To maintain clarity and focus, we chose to highlight the $\mathrm{GD}$$\mathrm{SGD}$ case, which already illustrates the core insights.
Many practical algorithms perform multiple local $\mathrm{SGD}$ steps to reduce communication overhead (e.g., $\mathrm{FedAvg}$). However, our focus in this work is not on practical communication efficiency, but rather on understanding the generalization behavior in the presence of misbehaving agents in the most direct way for simplicity and clarity. We thus chose to analyze the setting where each client performs a single gradient step which provides a clearer and more focused framework to present and understand the fundamental algorithmic stability properties. We believe our insights extend to the multi-step case, and see no reason why the general conclusion of our paper would be impacted. Exploring this formally is an interesting direction for future work.
Discussion on other threat model formulations.
Other threat models could be investigated in future work, building upon the proof techniques developed in our work. To offer some preliminary context, we briefly discuss two relevant threat model formulations below.
1. Locally-poisonous threat model. The recent work by Farhadkhani et al. (2024b), which studies the locally-poisonous threat model, indicates that this threat model tends to have a fundamentally lower impact on optimization error compared to the Byzantine failures and data poisoning threat models considered in our paper. Studying stability and generalization under locally-poisonous threat model is an interesting future work.
2. Backdoor (or targeted) attacks. While important, this threat model differs in goals and evaluation from our study of generalization error under untargeted attacks. That said, they can be seen as special cases of Byzantine failures and data poisoning. Extending our theory to handle targeted attacks like backdoors remains an interesting direction for future work.